A potential energy function for a system in which a two-dimensional force acts is of the form U=3x2y−4x. Find the x component of the force at point (2,3).
1
Expert's answer
2019-01-08T10:54:11-0500

The x-component of the force is

F_x=∂U/∂x=∂/∂x (3x^2 y-4x)=6xy-4

At this point:

F_x (2,3)=6(2)(3)-4=32 N.